Question 2 A. Suppose that the aggregate demand and supply curve for solar panels is given by P = 10 - 2Q and P = 1 + 5Q respectively. i. Draw the demand and supply curves for solar panels for this market and mark the equilibrium. Label this point as A. ii. Calculate the consumer surplus, producer surplus and total economic surplus of the market when it is in equilibrium. B. Suppose that, following a technological advancement, solar panels can be produced more cheaply. Furthermore, companies are purchasing more solar panels to improve their reputation. i. On a new graph, show and explain how the equilibrium price and quantity of solar panels will change due to both of these changes. Explain. ii. Graphically show the impact on consumer, producer and total economic surplus? Explain, why each of the “surplus’” have changed and why the previous equilibrium, A, is not Pareto efficient. In your answer you must show the Pareto improving transaction. C. Explain why the market equilibrium for a perfectly competitive market is also Pareto efficient. D. Assume that it is a perfectly competitive market. Explain why the long run the demand curve faced by the firm is horizontal. E. In the long run, how is the long-run supply curve derived? What is the shape of the supply curve and what does it represent? i.e. Is it upward sloping/downward sloping/horizontal/all of the aforementioned? Explain. In your answer, you must explain the assumptions which drive the shape of the curve and what would happen if these assumptions are violated.
Answer #1

(Part A)

(i)

From demand function, when Q = 0, P = 10 (vertical intercept) & when P = 0, Q = 10/2 = 5 (horizontal intercept).

From supply function, when Q = 0, P = 1 (vertical intercept)

In following graph, D0 and S0 are initial demand and supply curves, intersecting at point A with price P0 and quantity Q0.

(ii)

In equilibrium, demand = supply.

10 - 2Q = 1 + 5Q

7Q = 9

Q = 1.29

P = 10 - (2 x 1.29) = 10 - 2.58 = 7.42

Consumer surplus (CS) = (1/2) x (10 - 7.42) x 1.29 = (1/2) x 2.58 x 1.29 = 1.6641

Producer surplus (PS) = (1/2) x (7.42 - 1) x 1.29 = (1/2) x 6.42 x 1.29 = 4.1409

Total surplus (TS) = CS + PS = 1.6641 + 4.1409 = 5.805

NOTE: As HOMEWORKLIB Answering Policy, 1st part with multiple sub-parts has been answered.